OpenStudy (anonymous):

Indicate the equation of the given line in standard form. The line that is the perpendicular bisector of the segment whose endpoints are R(-1, 6) and S(5, 5)

OpenStudy (anonymous):

someone??

OpenStudy (bradely):

First find slope m= (y2-y1)/(x2-x1) =(5-6)/(5+1)=-1/6 Perpendicular slope m1 =-1/m =6 Midpoint of the points ,((x1+x2)/2,(y1+y2)/2) =((-1+5)/2,(6+5)/2) =(2,11/6) y-y1=m1(x-x1) (y-(11/6))=6(x-2) 6y-11 =36(x-2) 36x-6y-61=0 Source: http://www.mathskey.com/question2answer/ask